OpenStudy (anonymous):

at what point do the lines 5x +y=19 and-x +2y=6 intersect?

OpenStudy (anonymous):

Set them both equal to y so that you can then set them equal to each other. 5x+y=19 ---------> y=19-5x -x+2y= -6 --------->y= (x-6)/2 then substitute the y and set the equations equal to each other 19-5x= (x-6)/2 now solve for x 2(19-5x)=x-6 38-10x=x-6 44=11x x=4 plug 4 in for x in either one of the original equations 5(4)+y=19 y= -1 so the answer is (4, -1)

OpenStudy (anonymous):

thanks I understand now